OpenStudy (anonymous):

Need help. The equation of a function is y=2^x -4 The book wants me to find the coordinates of 2 given points, Which i have done. A = (0,-3) and B = (2,0). However the book is asking for the values of these points, in the answers it says A=4 and B=0.5 How do you do this? Please explain

OpenStudy (anonymous):

@Warlord , If possible please attach the original question because I am unable to understand your question

OpenStudy (anonymous):

the book seems to think that the coordinate (0,-3) on this graph made by y=2^x -4 has a value of 4. I am not sure how this works

OpenStudy (anonymous):

will a photo of the book work?

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

ok doing it now

OpenStudy (anonymous):

theres the photo

OpenStudy (paxpolaris):

it's asking for the co-ordinates for the Capital-letters (the points) and the values of the lower-case letters (the constants). A=(0,-3) is right a=4 is right

OpenStudy (anonymous):

umm im not sure what you mean by the values of the lower case letters, there are no lower case letters in this case

OpenStudy (paxpolaris):

\[\Large y=\sqrt{x+a}\]

OpenStudy (anonymous):

the function of the first graph is \[y=2^{x}-4\]

OpenStudy (anonymous):

so why is the value of the point (0,-3) four? and why is the value of the point (2,0) zero point five

OpenStudy (paxpolaris):

it is not. a= 4 is from the second equation in the list the points A,B,C,D,E,F,G,H,I have co-ordinates , not values.. the constants a,b,c,d,e,f,g,h,i have values

OpenStudy (anonymous):

OH dear god, i see now

OpenStudy (anonymous):

thankyou very much